How much will you spend if you spend $2.00 every 3 minutes for 9 days?
Answer by nerdybill(6962)

minutes in a day:
24 hrs * 60 min/hr = 1440 min per day
.
amount spent per day = 2 * (1440/3) = $960 per day
.
for nine days = 9*960 = $8,640